8, Calculate the pH of a solution that is 0.111 M in sodium formate (NaHCOO) and 0.253 Min formic acid (HCOOH). The pke of formic acid is 3.745 (a) 3.387 (b) 4.103 (c) 14.36 (d) 10.61 (e) 5.296 Use the following oxidation-reduction reaction to answer questions 9-11. Fe2+(aq)Mn04 (aq) Fe3+(aq)Mn2+(aq) 9. What are the coefficients in front of H and Fe in the balanced equation? (a) H 2, Fe3+ 3 (c) H. = 3, Fe3+ 2 10. Which chemical species is the strongest oxidizing ogent? (a) Fe(ag)
On Number 8 the answer is A but I don’t understand why
Answer #1

Answer:

Gievn buffer is formic acid and sodium formate.

pH=pKa+log[salt]/[acid]

Given pKa=3.745, [salt]=[NaHCOO]=0.111 M,

[Acid]=[HCOOH]=0.253 M.

Therefore pH=3.745 + log(0.111/0.253)

pH=3.745 - 0.357=3.387.

pH=3.387.

So option (A) is correct.

Please let me know if you have any doubt. Thanks and I hope you like it.